OpenStudy (thedj4jc):

Evaluate f-2g for f = 3 and g = 27. 3 9 This doesn't make sense... 3^2 = 9 and 9*27=243. So how could any of these be correct?

OpenStudy (anonymous):

3 - 2(27) = 3 - 54 = -51 , you sure you haven't made any mistakes ?

OpenStudy (thedj4jc):

It's f^-2g. I didn't see the - in it before.

OpenStudy (thedj4jc):

The answer's 3